# Language, Agent, and World Models

### v1

@@ Abstract
Meta titles and descriptions strongly shape engagement in search and recommendation platforms, yet optimizing them remains challenging. Search engine ranking models are black boxes environments, explicit labels are unavailable, and feedback such as click-through rate (CTR) arrives only post-deployment. Existing template, LLM, and retrieval-augmented approaches either lack diversity, hallucinate attributes, or ignore whether candidate phrasing has historically succeeded in ranking. This leaves a gap in directly leveraging implicit signals from observable outcomes. We introduce MetaSynth, a multi-agent retrieval-augmented generation framework that learns from implicit search feedback. MetaSynth builds an exemplar library from top-ranked results, generates candidate snippets conditioned on both product content and exemplars, and iteratively refines outputs via evaluator–generator loops that enforce relevance, promotional strength, and compliance. On both proprietary e-commerce data and the Amazon Reviews corpus, MetaSynth outperforms strong baselines across NDCG, MRR, and rank metrics. Large-scale A/B tests further demonstrate +10.26% CTR and +7.51% clicks. Beyond metadata, this work contributes a general paradigm for optimizing content in black-box systems using implicit signals.
1 Introduction
Search and recommendation systems are central to online discovery, yet their internal ranking functions are typically opaque covington2016deep ; doshi2017towards . These systems disclose little about how items are ordered, but their outputs such as search engine result pages (SERPs) consistently reflect stable preferences in the way information is phrased and structured. Among the most impactful elements are meta titles and descriptions, the short snippets displayed to users at decision time, which strongly influence click behavior and downstream trafficclarke2007influence . Optimizing these snippets represents a high-leverage intervention for organic acquisition. However, the optimization signal is fundamentally black-box: ranking models expose no gradients, and observable metrics such as impressions or click-through rate (CTR) are only available post-deployment, where exploration is expensive and feedback is biased by confounding factors such as position and popularity joachims2017unbiased ; islam2025correcting ; zhao2024mitigate .
@@ 1 Introduction
Existing strategies have notable limitations. Template-based generation provides consistency but lacks expressiveness and generalization across domains gatt2018survey ; chen2021generate . Prompt-only large language models (LLMs) generate fluent text but remain ungrounded, often hallucinating attributes or reverting to generic phrasing maragheh2025future . Retrieval-augmented generation (RAG) improves factual grounding, but retrieval is usually based solely on content similarity, disregarding whether candidate styles have historically been rewarded in ranking outcomes yu2024evaluation ; forouzandehmehr2025cal . As a result, current methods fail to fully exploit the implicit supervision embedded in black-box outputs cao2025writing .
We address this gap with MetaSynth, a model-based multi-agent, retrieval-augmented LM framework designed to “play the black-box search engine” by leveraging weak supervision from observable outcomes. MetaSynth constructs an exemplar success library by harvesting (query, metadata) pairs from top-ranked results, which we interpret as an implicit world model capturing ranking and click preferences. For a new webpage, an agentic retriever synthesizes plausible queries and retrieves relevant exemplars; when coverage is insufficient, it expands the library. A constrained LM generator proposes candidate snippets conditioned on content and exemplar styles. A panel of evaluator agents (critics) then scores relevance, coverage, promotional tone, and policy/brand compliance; a consensus coordinator integrates their feedback to plan targeted revisions. The result is an interpretable optimization loop guided by transparent objectives over a learned world model.
@@
  • We design an automated evaluator–generator refinement loop with consensus-driven feedback, ensuring outputs satisfy relevance, fluency, and brand/policy guardrails.
@@
  • We demonstrate MetaSynth’s effectiveness through both large-scale offline evaluations and online A/B tests, showing consistent improvements over strong baselines and significant real-world impact on CTR and clicks.
Figure 2: MetaSynth Framework: Three main modules to generate and optimize meta titles and descriptions for items, according to seller’s provided information, and constraints for better search engine ranking.
2 Related Work
@@
Search and recommendation systems expose only their outputs, making it difficult to infer or optimize internal ranking preferences rolinek2020optimizing ; sazanovich2021solving ; giahi2025systems . In this setting, user-facing snippets (meta titles and descriptions on SERPs) are known to shape attention and clicks, and thus downstream traffic chuklin2015anatomy ; islam2019micro . Prior work on query-biased summarization and snippet construction improves relevance and readability, but typically optimizes proxy text-quality metrics rather than outcome-driven objectives tied to ranking or clicks cao2016attsum ; baumel2018query ; chen2020abstractive . Moreover, exploration in production is expensive and biased by position and popularity, motivating methods that can learn from observational data without full access to the ranker joachims2017unbiased ; ermis2020learning .
@@
Template-based NLG offers control and consistency for product and listing metadata but struggles to generalize stylistically across domains reiter1995nlg ; deemter2005real ; gajbhiye2021template . Prompt-only LLMs increase fluency and diversity, yet can hallucinate attributes or regress to generic phrasing without grounding in historically successful styles chen2025carts ; maragheh2025future . Retrieval-augmented generation (RAG) improves faithfulness by conditioning on retrieved evidence, but standard retrieval is primarily similarity-driven and agnostic to whether candidate styles have performed well under ranking lewis2020retrieval ; maragheh2025arag ; forouzandehmehr2025cal . Recent work examines leveraging observational signals for writing and recommendation loghmani2025aligning ; cao2025writing ; chen2023bias , yet typically treats them as features for rerankers or classifiers rather than as priors for style-aware generation deldjoo2024recommendation ; khan2023learning .
@@
Multi-agent LLM frameworks and self-refinement protocols use specialized roles (e.g., critics and verifiers) to improve reliability and adherence to constraints yu2025table ; yuan2025reinforce ; chen2025carts ; wan2025mamm ; giahi2025vl . However, most evaluators optimize textual-quality proxies rather than outcome-aligned objectives and rarely close the loop with retrieval decisions motwani2024malt .
@@
In contrast to mentioned work MetaSynth is positioned at the intersection of RAG, weak supervision, and multi-agent self-improvement. It builds an exemplar library from top-ranked results and conditions generation jointly on product content and outcome-informed stylistic exemplars, thereby injecting an explicit, outcome-aligned prior into the generator rather than relying on content similarity alone. An agentic retriever synthesizes queries to retrieve and expand coverage when library support is sparse, while a constrained generator produces candidates that are subsequently refined by a panel of evaluator agents scoring relevance, coverage, promotional strength, and brand/policy compliance; a consensus coordinator converts this feedback into targeted revisions. This closes the loop between retrieval, generation, and evaluation without gradient access to the ranker or curated preference labels, yielding an interpretable and controllable optimization process that aligns style with observed preferences. As such, MetaSynth operationalizes weak supervision for style-aware generation and offers a practical pathway to optimize text for engagement in black-box ranking environments.
3 Methodology
@@
3.1 Problem Definition
Let $\mathcal{X}$ denote the set of retailer product pages for a target eCommerce platform. For a page $x\in\mathcal{X}$, with its associated textual meta data like product name, brand, category, seller description $\mathbf{a}(x)$, the goal is to produce a meta-snippet $y=(\tau,\delta)$ comprising a meta title $\tau$ and meta description $\delta$ that maximizes downstream organic acquisition under a black-box search engine. We denote the (unknown) objective induced by the search engine and user behavior as
@@ 3.1 Problem Definition
3.2 Library Construction Module: Offline Vector DB Construction
Given a seed set of popular queries $\mathcal{Q}_{\mathrm{S}}$, the Library Construction agent issues calls to the black-box engine and builds $\mathcal{L}=\bigcup_{q\in\mathcal{Q}_{\mathrm{S}}}\,\{(q,u_{i},\tau_{i},\delta_{i},r_{i})\}_{i=1}^{K_{\mathrm{lib}}}.$ For saving each exemplar $i$, both its meta title $\tau_{i}$ and meta description $\delta$ are embedded using $\mathbf{g}_{y_{i}}(e_{i})=\mathbf{g}_{y_{i}}(\tau_{i}\!\oplus\!\delta_{i})$. We do not save the embedding vector if the cosine similarity with the most similar item in the library is more than a threshold $\epsilon_{\mathrm{dup}}$. In other words, $e_{i}$ is considered a duplicate of existing $e_{j}$ in the library if : $\mathrm{sim}(\mathbf{g}_{y}(e_{i}),\mathbf{g}_{y}(e_{j}))>\epsilon_{\mathrm{dup}}$.
@@ 3.2 Library Construction Module: Offline Vector DB Construction
3.3 Generation Module: Generation via Agentic Retrieval-Augmented Few-Shot Prompting
3.3.1 Target Query Detection and Library Construction
@@
and then invokes $\mathsf{Search}$ for each $q_{\mathrm{new},x}$.
@@
3.3.2 Exemplar set construction and Generation
Let $\mathcal{F}_{x}$ be the candidate exemplars set for webpage $x$. The $m$ few-shot exemplars selected by a greedy algorithm using Maximal Marginal Relevance (MMR) carbonell1998use to balance relevance and diversity:
@@ 3.3.2 Exemplar set construction and Generation
where $\mathcal{F}$ is the selected growing set and $\lambda\in[0,1]$. We iterate $m$ steps to obtain $\mathcal{F}_{x}$. This will ensure that we have a diverse enough set of successful examples to be passed to meta-snippet generation.
$y^{(0)}\;=\;G\big(x,\mathcal{F}_{x},\mathcal{B}\big),$ (6)
3.4 Evaluation Module: Evaluation Agents and Iterative Refinement
In this case, $\mathbf{c}^{(t)}$ is injected as structured constraints (e.g., “increase promotional strength,” “insert CTA,” “remove forbidden term hh”). The cycle stops at iteration $t^{\star}$ if either (i) if the Evaluator accepts the generated text on all criteria or (ii) iteration hits the max iteration budget $K_{\max}$. Optionally, a stagnation rule halts the iterations if enough improvement does not happen at all (or enough) for two consecutive steps. Fig. 3 shows an example of how evaluation module modifies the initial generated output of 3.3.2. Please refer to Appendix B for entire algorithm.
4 Experiments and Results
4.1 Data and Experiment Setting
  • Amazon Review dataset hou2024bridging : A widely used public benchmark curated by McAuley Lab. We extract 30,000 items across three domains, Home, Toys, and Fashion, with 10,000 items in each category. This dataset provides structured metadata such as item descriptions, prices, and review text, making it complementary to our proprietary corpus.
For preprocessing, we apply standard NLP techniques to remove non-ASCII characters, special symbols, and noise. For each item, we employ an LLM to generate the most likely user query that could lead to that item. We then identify the closest matching query from our VectorDB (as described in section 3.3.1) and retrieve the top-$k$ exemplar meta titles and descriptions. These exemplars serve as weakly supervised demonstrations for the Generation Module (section 3.3.2), ensuring that outputs are conditioned on successful historical styles. We further regenerate meta-snippets based on feedback and use them as outputs to compare again these three benchmarks (section 3.4). This setup allows us to test MetaSynth on both controlled proprietary environments and an open, publicly reproducible benchmark.
4.2 Evaluation Metrics
We compare MetaSynth against three representative baselines:
  1. 1.
  2. 2.
    DRE (Direct Retrieval Enhancement) gao2024dre : An approach that extracts keywords from items and incorporates them into generation prompts, thereby enriching outputs with content-derived terms.
  3. 3.
    CoT (Chain-of-Thought prompting) wei2022chain : A reasoning-based method that guides LLMs through structured intermediate steps, improving factual alignment between metadata and generated snippets.
To ensure consistent evaluation, we use GPT-4.1-mini as a judgment model, applying the LLM-as-a-judge paradigm zheng2024mtbench to rank snippets generated by each of these baseline models. We report three widely accepted ranking metrics:
4.3 Experiment Results
Table 1: Offline evaluation results: LLM-as-a-judge metrics across Amazon and proprietary datasets for meta titles and meta descriptions.
Dataset Approach Meta Title Meta Description
NDCG MRR Avg Rank NDCG MRR Avg Rank
Amazon Fashion Baseline 0.5970 0.4617 2.5284 0.5504 0.4013 2.8355
DRE 0.5095 0.3505 3.2731 0.5673 0.4261 2.9335
COT 0.6280 0.5046 2.4911 0.6443 0.5282 2.5221
MetaSynth 0.8190 0.7601 1.7025 0.7911 0.7213 1.6987
Amazon Toys Baseline 0.6302 0.5051 2.3450 0.5873 0.4492 2.5834
DRE 0.5174 0.3606 3.2197 0.5972 0.4657 2.7714
COT 0.5911 0.4562 2.7260 0.5804 0.4450 2.9459
MetaSynth 0.8149 0.7551 1.7015 0.7881 0.7169 1.6921
Amazon Home Baseline 0.6648 0.5510 2.2187 0.5609 0.4147 2.7284
DRE 0.5486 0.4005 3.0045 0.5885 0.4543 2.8140
COT 0.5846 0.4478 2.8052 0.6048 0.4764 2.7941
MetaSynth 0.7598 0.6809 1.9694 0.7996 0.7316 1.6627
Proprietary Baseline 0.5762 0.4352 2.6834 0.4771 0.3160 3.2655
DRE 0.5553 0.4102 2.9622 0.5066 0.3552 3.1309
COT 0.6527 0.5381 2.4191 0.7117 0.6243 1.9640
MetaSynth 0.7631 0.6882 1.9716 0.7835 0.7204 1.6416
On the Amazon Review dataset (See Table 1) , Baseline establishes a modest benchmark (e.g., NDCG 0.5970, MRR 0.4617 on Fashion titles), but its lack of explicit reasoning often results in suboptimal ranking. DRE provides incremental gains in some settings, such as improved meta description retrieval on Home (NDCG 0.5885 vs. 0.5609 for Baseline), but overall lags behind other methods, as its heuristic adjustments fail to capture deeper semantic structure. COT delivers more consistent improvements by incorporating structured reasoning, achieving higher retrieval quality (e.g., Fashion meta description MRR of 0.5282 vs. 0.4013 for Baseline). However, while COT narrows the gap, its reasoning alone cannot fully address domain-specific nuances in product data.
MetaSynth achieves the strongest results across all datasets, significantly outperforming prior methods. On Fashion titles, MetaSynth reaches NDCG of 0.8190 and MRR of 0.7601, compared to 0.6280 and 0.5046 for COT. A similar trend holds for meta descriptions, where MetaSynth (NDCG 0.7911, MRR 0.7213) consistently outperforms all baselines. The improvements extend to Toys (title MRR 0.7551 vs. 0.4562 for COT) and Home (description NDCG 0.7996 vs. 0.6048 for COT), with MetaSynth reducing average rank to  1.7 across domains. These results highlight MetaSynth’s ability to generalize across categories while preserving fine-grained detail in retrieval.
The offline evaluation metrics clearly show that our proposed approach MetaSynth consistently performs in all of them beating all benchmark models indicating the importance of exemplar library and evaluator feedback loop. Although benchmarks like COT and DRE fare better than Vanilla, MetaSynth yields state-of-the-art retrieval performance across all datasets and metrics. Please refer to Appendix A for real examples from our experiments showing the results of each benchmark model.
4.4 Ablation Study
Table 2: Ablation Studies
Approach Average rank NDCG MRR
MetaSynth wo RAG 2.4830 0.6245 0.5018
Meta Synth wo Evaluation 1.9769 0.7267 0.6353
MetaSynth 1.6416 0.7835 0.7204
To systematically analyze the contribution of individual components within our pipeline, we conducted ablation studies across 3 variants i) MetaSynth without Library Construction Module ii) MetaSynth without Evaluation module iii) Complete MetaSynth pipeline. For this analysis, we compared the performance of each these variants with benchmark models on our Proprietary dataset.
The results summarized in (see Table LABEL:Table:Ablation ) show that exclusion of Library Construction Module leads to a substantial degradation in performance there is a 33% drop in average rank compared to the complete Meta Synth framework which highlights the role of high quality exemplar titles and descriptions. Additionally, when we remove the Evaluation and Feedback module, there is again a considerable decline in all metrics indicating that the loop of evaluation, feedback consolidation and regeneration is essential for good performance. In ranking effectiveness, MetaSynth yields a 25.5% gain in MRR and a 25.4% gain in NDCG compared to MetaSynth w/o RAG, while achieving a further 13.4% (MRR) and 7.8% (NDCG) improvement over MetaSynth w/o Evaluation. Notably, the observed performance drop is more pronounced when Library Module is omitted compared to ii) W/o Evaluation suggesting that Library Module exerts a greater influence on the performance, Overall, having all three components surpasses the other methods demonstrating the significance of each of the components. In particular, evaluation modules enhances ranking precision, while Library Construction Module enriches the VectorDB enabling Meta Synth to perform the best in all offline evaluation metrics.
4.5 A/B Test Results
To evaluate the impact of meta-titles and snippets on user engagement, we conducted 4-weeks long A/B test to compare a propreitary control search engine meta generator with the ones generated through MetaSynth. MetaSynth leads to a +10.26% improvement in clicks and 7.51% in CTR when compared to the control model (Table 3). These statistically significant lifts confirm that MetaSynth’s offline improvements translate directly into real-world user engagement. Importantly, the gains substantially outweigh the additional inference costs introduced by the multi-agent pipeline, demonstrating that MetaSynth is both effective and scalable for large catalogs.
Table 3: A/B test performance
Metric Lift
CTR +10.26%
Overall Clicks +7.51%
5 Conclusion
Beyond metadata generation, this work contributes a broader paradigm for learning from implicit feedback in black-box environments. By showing how multi-agent generation can integrate weak supervision, retrieval, and iterative critique, we open new directions for applying similar techniques to recommendation, personalization, and ranking-adjacent tasks. Future research may extend this framework to richer modalities (e.g., images, video), integrate counterfactual debiasing of implicit signals, and explore theoretical guarantees for convergence under noisy supervision.
Appendix A Case studies
Figure 4: Case studies comparing MetaSynth method with three other studies (baseline, DRE, COT), in which MetaSynth outcome ranked first among others in the search engine. Some influential words that might be affected the ranking are demonstrated with bold font.
Below we compare the four variants for the “US Capitol Building and Red Balloons” canvas (See Figure 4, top-left case). Relative to the Baseline, DRE, and CoT texts, the MetaSynth description exhibits stronger lexical economy and discourse naturalness: it foregrounds the core entity and creator (“…by Fab Funky”), compresses categorical qualifiers (“modern & contemporary style”), and uses a single, semantically rich construction (“gallery‑wrapped masterpiece”) rather than a list of loosely coupled attributes. In contrast, Baseline omits brand/creator and leans on generic language (“brighten your walls”), DRE reads as a keyword list with rigid slot filling (“Modern, gallery‑wrapped, 24x32 inches”), and CoT over‑enumerates proper nouns and media terms (“Trademark Fine Art … Canvas Art … Giclée Print”) in a way that resembles inventory metadata rather than user‑centric copy. This shift from enumeration to fluent phrasing improves readability while preserving key tokens that matter for retrieval (e.g., “gallery‑wrapped”, style cues), thus aligning with relevance, readability, and technical‑compliance components of the surrogate objective. Second, the MetaSynth variant better harmonizes intent coverage with precision by balancing audience framing and attribute salience. The phrase “modern & contemporary style” broadens matchability across adjacent intents without resorting to repetitive keyword stuffing, while mentioning the creator “Fab Funky” supplies a trusted brand cue that can disambiguate entity searches. Compared with DRE’s mechanical cadence and CoT’s concatenated title‑like string, MetaSynth’s clause structure distributes salient tokens across sentences (lead with identity $\rightarrow$ situate style $\rightarrow$ state mounting/finish) to satisfy evaluator checks on relevance and detail without triggering duplication penalties.
Additionally, we compare the four variants for the “Walker Edison Farmhouse Sliding Door Cabinet" product description (See Figure 4, top-right case). Relative to the Baseline, DRE, and CoT texts, the MetaSynth description demonstrates superior lexical efficiency: it leads with action-oriented framing (“Organize in style"), emphasizes core value propositions (ample storage space, durable materials), and maintains brand recognition while avoiding specification overload. In contrast, Baseline and DRE rely on identical phrasing, while CoT overwhelms with technical specifications (68"H x 36"L, “grooved sliding door") that read as catalog entries rather than persuasive copy. MetaSynth’s selective emphasis highlights functionality and longevity without dimensional clutter, while “Organize in style" efficiently combines utility with aesthetic appeal. This creates broader search matching than CoT’s specification-heavy approach while maintaining more substance than Baseline’s generic “rustic charm" language, improving both readability and conversion potential.
Third, we compare the four variants for the “Avia Women’s Transition V-Neck T-Shirt" product description (See Figure 4, bottom-left). Relative to the Baseline, DRE, and CoT texts, the MetaSynth description demonstrates superior market positioning: it leads with aspirational framing (“Upgrade your activewear collection"), strategically emphasizes category context (activewear) and tactile appeal (soft hand feel), while maintaining technical benefits within engaging copy. In contrast, Baseline and DRE open with generic comfort claims, while CoT defaults to feature enumeration (“Moisture Wicking Fabric & Soft Peached Jersey. Relaxed Fit") that prioritizes specifications over lifestyle integration. MetaSynth’s selective bolding of activewear establishes clear category positioning for broader intent matching, while soft hand feel translates technical “peached jersey" into consumer-friendly sensory language. The phrase “Upgrade your activewear collection" positions the product as enhancement rather than necessity, creating stronger purchase motivation than CoT’s functional listing or Baseline’s passive “stay cool" messaging, improving both category relevance and conversion appeal.
Finally, we compare the four variants for the “NHL Portable Bar featuring Buffalo Sabres" product description (See Figure 4, bottom-right). Relative to the Baseline, DRE, and CoT texts, the MetaSynth description demonstrates superior contextual targeting: it leads with experiential framing (“Get ready to elevate your next picnic"), emphasizes specific use scenarios (your next picnic) and key functional benefits (spacious), while streamlining technical details into essential selling points. In contrast, Baseline opens with generic “game day" positioning, DRE focuses heavily on structural specifications, while CoT overwhelms with dimensional data (39" x 15") and feature lists that prioritize inventory details over lifestyle appeal. MetaSynth’s selective bolding of your next picnic creates direct personal relevance and expands beyond traditional sports contexts, while spacious translates technical shelf configurations into practical consumer benefit language. The phrase “Get ready to elevate your next picnic" positions the product within broader outdoor entertainment rather than limiting to sports events, creating wider intent matching than Baseline’s “game day" restriction or CoT’s specification-heavy approach, improving both contextual relevance and purchase motivation.
Appendix B Algorithm
In this section of the appendix, we present a pseudocode for the proposed methodology of MetaSynth.

### v2

# MetaSynth: Multi-Agent Metadata Generation from Implicit Feedback in Black-Box Systems

@@ Abstract
Product titles and descriptions displayed in search results (also called meta title and descriptions) strongly shape engagement in search and recommendation platforms, yet optimizing them remains challenging. Ranking models are black boxes, explicit labels are unavailable, and feedback such as click-through rate (CTR) arrives only post-deployment. Existing template, LLM, and retrieval-augmented approaches either lack diversity, hallucinate attributes, or ignore whether candidate phrasing has historically succeeded in ranking. This leaves a gap in directly leveraging implicit signals from observable outcomes. We introduce MetaSynth, a multi-agent retrieval-augmented generation framework that learns from implicit search feedback. MetaSynth builds an exemplar library from top-ranked results, generates candidate snippets conditioned on both product content and exemplars, and iteratively refines outputs via evaluator–generator loops that enforce relevance, promotional strength, and compliance. On both proprietary e-commerce data and the Amazon Reviews corpus, MetaSynth outperforms strong baselines across NDCG, MRR, and rank metrics. Large-scale A/B tests further demonstrate +10.26% CTR and +7.51% clicks. Beyond metadata, this work contributes a general paradigm for optimizing content in black-box systems using implicit signals.
I Introduction
Search and recommendation systems are central to online discovery, yet their internal ranking functions are typically opaque [covington2016deep, doshi2017towards]. These systems disclose little about how items are ordered, but their outputs such as search engine result pages (SERPs) consistently reflect stable preferences in the way information is phrased and structured. Among the most impactful elements are meta titles and descriptions, the short snippets displayed to users at decision time, which strongly influence click behavior and downstream traffic[clarke2007influence]. Optimizing these snippets represents a high-leverage intervention for organic acquisition. However, the optimization signal is fundamentally black-box: ranking models expose no gradients, and observable metrics such as impressions or click-through rate (CTR) are only available post-deployment, where exploration is expensive and feedback is biased by confounding factors such as position and popularity [joachims2017unbiased, islam2025correcting, zhao2024mitigate].
@@ I Introduction
Existing strategies have notable limitations. Template-based generation provides consistency but lacks expressiveness and generalization across domains [gatt2018survey, chen2021generate]. Prompt-only large language models (LLMs) generate fluent text but remain ungrounded, often hallucinating attributes or reverting to generic phrasing [maragheh2025future]. Retrieval-augmented generation (RAG) improves factual grounding, but retrieval is usually based solely on content similarity, disregarding whether candidate styles have historically been rewarded in ranking outcomes [yu2024evaluation, forouzandehmehr2025cal]. As a result, current methods fail to fully exploit the implicit supervision embedded in black-box outputs [cao2025writing].
We address this gap with MetaSynth, a multi-agent retrieval-augmented generation framework designed to “play the black box search engine” by systematically leveraging weak supervision from observable outcomes. MetaSynth constructs a “exemplar library” by harvesting (query, metadata) pairs from top-ranked results, thereby encoding implicit ranking preferences into a reusable corpus. For a new webpage and for constricting the search engine meta snippets of the webpage, an agentic retriever synthesizes plausible queries, retrieves the relevant examples from the library, and enriches the exemplar library when there is no relevant example in it. A constrained generator then produces candidate snippets conditioned on both product content and exemplar styles. These candidates are iteratively evaluated by a panel of specialized evaluator agents that assess relevance, coverage, promotional tone, and compliance with brand constraints. Their feedback is fused by a consensus coordinator into targeted revisions, yielding an interpretable optimization loop guided by transparent objectives.
@@
  • We design an automated evaluator–generator refinement loop with consensus-driven feedback, ensuring that outputs satisfy relevance, fluency, and brand/policy guardrails.
@@
  • We demonstrate the effectiveness of MetaSynth through both large-scale offline evaluations and online A/B tests, showing consistent improvements over strong baselines and significant real-world impact on CTR and clicks.
Figure 2: MetaSynth Framework: Three main modules to generate and optimize meta titles and descriptions for items, according to seller’s provided information, and constraints for better search engine ranking. Left block represents the RAG-based exemplar library creation. Middle block shows the generation module for both meta title and description based on seller provided data and exemplar library. Right block represents the evaluation/refinement loop, where multi evaluator agents evaluate the title/description for their assigned criteria, and if failed to satisfy that criteria, the feedback consolidation agent directs such title/description to re-generation agent along with the provided feedback from rejecting evaluator.
II Related Work
@@
Search and recommendation systems expose only their outputs, making it difficult to infer or optimize internal ranking preferences [rolinek2020optimizing, sazanovich2021solving, giahi2025systems]. In this setting, user-facing snippets (meta titles and descriptions in SERPs) are known to shape attention and clicks, and thus downstream traffic [chuklin2015anatomy, islam2019micro]. Prior work on query-biased summarization and snippet construction improves relevance and readability, but typically optimizes proxy text-quality metrics rather than outcome-driven objectives tied to ranking or clicks [cao2016attsum, baumel2018query, chen2020abstractive]. Moreover, exploration in production is expensive and biased by position and popularity, motivating methods that can learn from observational data without full access to the ranker [joachims2017unbiased, ermis2020learning].
@@
Template-based NLG offers control and consistency for product and listing metadata but struggles to generalize stylistically across domains [reiter1995nlg, deemter2005real, gajbhiye2021template]. Prompt-only LLMs increase fluency and diversity, yet can hallucinate attributes or regress to generic phrasing without grounding in historically successful styles [chen2025carts, maragheh2025future]. Retrieval-augmented generation (RAG) improves faithfulness by conditioning on retrieved evidence, but standard retrieval is primarily similarity-driven and agnostic to whether candidate styles have performed well under ranking [lewis2020retrieval, maragheh2025arag, forouzandehmehr2025cal]. Recent work examines leveraging observational signals for writing and recommendation [loghmani2025aligning, cao2025writing, chen2023bias], yet typically treats them as features for rerankers or classifiers rather than as priors for style-aware generation [deldjoo2024recommendation, khan2023learning].
@@
Multi-agent LLM frameworks and self-refinement protocols use specialized roles (e.g., critics and verifiers) to improve reliability and adherence to constraints [yu2025table, yuan2025reinforce, chen2025carts, wan2025mamm, giahi2025vl]. However, most evaluators optimize textual-quality proxies rather than outcome-aligned objectives and rarely close the loop with retrieval decisions [motwani2024malt].
@@
In contrast to mentioned work, MetaSynth treats the SERP itself as weak supervision about which phrasing and structure succeed under a black-box policy. It builds an exemplar library from top-ranked results and conditions generation jointly on product content and outcome-informed stylistic exemplars, thereby injecting an explicit, outcome-aligned prior into the generator rather than relying on content similarity alone. An agentic retriever synthesizes queries to retrieve and expand coverage when library support is sparse, while a constrained generator produces candidates that are subsequently refined by a panel of evaluator agents scoring relevance, coverage, promotional strength, and brand/policy compliance; a consensus coordinator converts this feedback into targeted revisions. This closes the loop between retrieval, generation, and evaluation without gradient access to the ranker or curated preference labels, yielding an interpretable and controllable optimization process that aligns style with observed preferences. As such, MetaSynth operationalizes weak supervision for style-aware generation and offers a practical pathway to optimize text for engagement in black-box ranking environments.
III Methodology
@@
III-A Problem Definition
Let $\mathcal{X}$ denote the set of retailer product pages for a target eCommerce platform. For a page $x\in\mathcal{X}$, with its associated textual meta data like product name, brand, category, seller description $\mathbf{a}(x)$, the goal is to produce a meta-snippet $y=(\tau,\delta)$ comprising a meta title $\tau$ and meta description $\delta$ that maximizes downstream organic acquisition from a black-box search engine. We denote the (unknown) objective induced by the search engine and user behavior as
@@ III-A Problem Definition
III-B Library Construction Module: Offline Vector DB Construction
Given a seed set of popular queries $\mathcal{Q}_{\mathrm{S}}$, the Library Construction agent issues calls to the black-box engine and builds $\mathcal{L}=\bigcup_{q\in\mathcal{Q}_{\mathrm{S}}}\,\{(q,u_{i},\tau_{i},\delta_{i},r_{i})\}_{i=1}^{K_{\mathrm{lib}}}.$ For saving each exemplar $i$, both its meta title $\tau_{i}$ and meta description $\delta$ are concatenated (denoted by $\oplus$) and embedded using $\mathbf{g}_{y_{i}}(e_{i})=\mathbf{g}_{y_{i}}(\tau_{i}\!\oplus\!\delta_{i})$. We do not save the embedding vector if the cosine similarity with the most similar item in the library is more than a threshold $\epsilon_{\mathrm{dup}}$. In other words, $e_{i}$ is considered a duplicate of existing $e_{j}$ in the library if : $\mathrm{sim}(\mathbf{g}_{y}(e_{i}),\mathbf{g}_{y}(e_{j}))>\epsilon_{\mathrm{dup}}$.
@@ III-B Library Construction Module: Offline Vector DB Construction
III-C Generation Module: Generation via Agentic Retrieval-Augmented Few-Shot Prompting
III-C1 Target Query Detection and Library Construction
@@
and then invokes $\mathsf{Search}$ for each $q_{\mathrm{new},x}$. This will ensure exemplar writings for all rounds of writing.
@@
III-C2 Exemplar set construction
Let $\mathcal{F}_{x}$ be the candidate exemplars set for webpage $x$. The $m$ few-shot exemplars selected by a greedy algorithm using Maximal Marginal Relevance (MMR) [carbonell1998use] to balance relevance and diversity:
@@ III-C2 Exemplar set construction
where $\mathcal{F}$ is the selected growing set and $\lambda\in[0,1]$. We iterate $m$ steps to obtain $\mathcal{F}_{x}$. This will ensure that we have a diverse enough set of successful examples to be passed to meta-snippet generation. Note that when selecting the exemplar set, we can potentially inflate the similarity score ($\lambda\cdot\mathrm{sim}\!\big(\mathbf{z}_{x},\mathbf{g}_{y}(e)\big)$) as a function of ranking, $r$, of the exemplar writing recorded in the exemplar library. In this way, we also prioritize writings with better rankings. One can potentially deflate too in cases where popularity bias is severely inherent in search engine’s recommendations.